# The Dynamic, the Static, and the Weak factor models and the analysis of high-dimensional time series

### v1

@@ 7 Undetected factors: “don’t throw out the baby with the bathwater!”
Now, undetected factors remain undetected because their empirical finite-$(n,T)$ cross-correlations are “small.” The benefits of detecting such factors and incorporating them into the common component is that their cross-correlations, there, would be exploited—which they are not in a componentwise or sparse analysis of the idiosyncratic. Since, however, these cross-correlations are too small to be detected, these benefits are small, too: leaving them in the idiosyncratic is basically harmless—as long as the idiosyncratic is not discarded!
The problem of undetected factors, thus, perhaps, is not as crucial, in practice, as it seems: if a factor gets undetected, the benefits of correctly identifying it as a factor are negligible.
@@ 8 Conclusions
    Since, as shown by Gersing et al. (2023), it is nesting the classical and widespread static model of Chamberlain (1983) and Chamberlain and Rothschild (1983), the General Dynamic Factor Model of Forni et al. (2000) is uniformly preferable—the benefit being the incorporation into the common space of the GRD-weak factors which, in the static approach, are treated as idiosyncratic;
@@
    Gersing-Rust-Deistler-weak factors “are everywhere,” but are incorporated into the space of common components in the GDFM approach—which explains the empirical finding that the dynamic approach outperforms the static one even when the assumptions of the static factor model are satisfied;
@@
    factor models are not a dimension-reduction technique and idiosyncratic components (which are not error terms) should not be discarded, as they may be large and have high predictive power;
  5. (v)
    in particular, estimated idiosyncratic components may contain undetected rate-strong factors: provided that idiosyncratic components are not discarded, these undetected factors are not lost, though.
These conclusions are likely to extend to matrix- and tensor-valued factor models, and to spatio-temporal ones, for which, however, a dynamic approach similar to the GDFM still needs to be developed.
Appendix. The static approximate factor model revisited.
In this appendix, we formally establish the fact that the representation result (4.2), along with assumption (4.3) on the eigenvalues $\lambda^{(n)}_{X;j}$, $j=1,\ldots,n$, $n\in{\mathbb{N}}$ of the covariance matrix ${\bm{\Gamma}}^{(n)}_{X}\coloneqq{\rm E}[{\bf X}^{(n)}_{t}{\bf X}^{(n)\prime}_{t}]$ of ${\bf X}^{(n)}_{t}\coloneqq(X_{1t},\ldots,X_{nt})^{\prime}$ imply that the assumptions (a)–(g) of the static approximate factor model (2.1) of Chamberlain and Rothschild are satisfied. Essentially, this consists in showing that the covariance matrices ${\bm{\Gamma}}^{(n)}_{\chi}\coloneqq{\rm E}[{\bm{\chi}}^{(n)}_{t}{\bm{\chi}}^{(n)\prime}_{t}]$ and ${\bm{\Gamma}}^{(n)}_{\xi}\coloneqq{\rm E}[{\bm{\xi}}^{(n)}_{t}{\bm{\xi}}^{(n)\prime}_{t}]$ of the statically common and idiosyncratic components ${\bm{\chi}}^{(n)}_{t}\coloneqq(\chi_{1t}^{\scriptscriptstyle\text{\rm stat}},\ldots,\chi_{nt}^{\scriptscriptstyle\text{\rm stat}})^{\prime}$ and ${\bm{\xi}}^{(n)}_{t}\coloneqq(\xi_{1t}^{\scriptscriptstyle\text{\rm stat}},\ldots,\xi_{nt}^{\scriptscriptstyle\text{\rm stat}})^{\prime}$ defined in (4.2) and their eigenvalues are such that
@@
    the $r$ eigenvalues $\lambda^{(n)}_{\chi;1},\ldots,\lambda^{(n)}_{\chi;r}$ of ${\bm{\Gamma}}^{(n)}_{\chi}$ tend to infinity as $n\to\infty$ (it readily follows from (i) that $\lambda^{(n)}_{\chi;j}=0$ for all $n\geq j>r$);
@@
${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}=\left\{z=\sigma\lim_{n\to\infty}w^{(n)}_{*}{\big{|}}\,(w^{(n)}_{*})_{n\in{\mathbb{N}}}\in{\mathcal{W}}_{r\rceil*},\,\sigma\in{\mathbb{R}}_{+}\right\}.$

### v3

@@ 8 Identifying the number of factors
One may wonder, however, whether this is worth the effort. Identifying hardly detectable factors and incorporating them in the common space, indeed, has limited practical impact: undetected factors remain undetected because their empirical finite-$(n,T)$ cross-correlations are “small.” The benefits of detecting such factors and incorporating them into the common component is that their cross-correlations, there, would be fully exploited—which they are not in a componentwise (no cross-correlation exploited) or sparse (only a few cross-correlations exploited) analysis of the idiosyncratic. Since these cross-correlations are too small to be detected, these potential benefits are small, too, and failing to detect these factors is basically harmless. The problem of undetected factors, thus, is perhaps not as crucial, in practice, as it seems—provided, however, that factor models are not considered as a dimension-reduction technique throwing away idiosyncratic components—an attitude that is adopted by several authors.
Whether static or dynamic, undetected factors, indeed, are not lost, but wrongly labeled as idiosyncratic. In a dimension-reduction perspective, the common component $\chi_{it}$ (be it $\chi_{it}^{\scriptscriptstyle{\text{\rm stat}}}$ or $\chi_{it}^{\scriptscriptstyle{\text{\rm dyn}}}$) is considered an approximation of the high-dimensional observation $X_{it}$ which, being reduced rank, is more tractable than $X_{it}$ itself. The idiosyncratic $\xi_{it}$ (either $\xi_{it}^{\scriptscriptstyle{\text{\rm stat}}}$ or $\xi_{it}^{\scriptscriptstyle{\text{\rm dyn}}}$) then is discarded as if it were a negligible error term—a somewhat regrettable (errors, in principle, are memoryless, hence cannot be predicted) terminology used by some authors and originating in the exact factor model literature where, by definition, no idiosyncratic lagged cross-correlation is present. Depending on the objective of the analysis or the nature of the dataset, this may or may not be legitimate. In an analysis aiming at constructing a few indices that summarize the “global state” of the economy and its impact on individual cross-sectional items, for instance, extracting the factors and projecting observed $X_{it}$’s on the factor space (yielding the common component $\chi_{it}$) is the obvious way to go: a dimension-reduction attitude, thus, in which idiosyncratic components safely can be ignored. The same attitude is common practice, e.g., in macroeconometric forecasting, where it seems that the idiosyncratic safely can be neglected so that forecasting $X_{i,t+h}$ boils down to forecasting the value at time $t+h$ of the common component, be it static or dynamic (Boivin and Ng, 2006; Luciani, 2014).
Whether static or dynamic, idiosyncratic components need not be small, though, and may be strongly autocorrelated, hence enjoying high predictive value for $X_{i,t+h}$. Discarding them, in a prediction context, is potentially quite damaging. This is often the case in financial applications dealing with panels of volatility measures (Herskovic et al., 2016; Barigozzi and Hallin, 2017, 2020). Rather than a dimension-reduction technique, factor models then should be considered a “divide and conquer” procedure where the common and the idiosyncratic are analyzed via distinct appropriate methods—reduced-rank-process techniques for the common, taking advantage of strong and pervasive cross-correlations, and componentwise or sparse techniques for the idiosyncratic, where little or no advantage can be gained from the negligible potential cross-correlations.
In forecasting problems, depending on the type (static or dynamic) of factor model adopted, this “divide and conquer” approach yields $h$-step ahead forecasts ${\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$ (based on the past $\{\chi_{js}^{\scriptscriptstyle{\text{\rm stat}}}|\,j\in{\mathbb{N}},s\leq\leavevmode\nobreak\ \!t\}$of all common components) of $\chi_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$ and ${\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$ (based, e.g., on the past $\{\xi_{is}^{\scriptscriptstyle\text{\rm stat}}|\,s\leq t\}$ of the $i$th idiosyncratic component only) of $\xi_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$, or forecasts ${\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}$ (based on $\{\chi_{js}^{\scriptscriptstyle{\text{\rm dyn}}}|\,j\in{\mathbb{N}},s\leq t\}$) of $\chi_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$ and ${\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}$ (based on $\{\xi_{is}^{\scriptscriptstyle\text{\rm sdyn}}|\,s\leq t\}$) of $\xi_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$. In these examples, a componentwise univariate forecasting strategy is considered for idiosyncratic components. More efficient idiosyncratic prediction strategies could be adopted, where several past idiosyncratic values are taken into account (as in sparse VAR prediction methods); in theory, one even could use all past idiosyncratic values, which in practice is rapidly infeasible, though. Mutatis mutandis, the results below would remain unchanged. For simplicity of exposition and notation, we therefore restrict our predictors of ${\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$ and ${\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$ to the componentwise forecasts $\xi_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$ just described.
After being obtained separately, however, these forecasts somehow should be brought back together to produce a forecast of $X_{i,t+h}$. To this end, one naturally could consider their sums
${X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}\coloneqq{\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}+{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$ (9.1)
and
${X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}\coloneqq{\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}+{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}.$ (9.2)
as predictors for $X_{i,t+h}$. This is what Fan et al. (2023) are doing in the context of static approximate factor models and Barigozzi et al. (2024) in the GDFM context, with sparse VAR empirical forecasts $\widehat{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$ and $\widehat{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}$ of ${\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$ and ${\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$, respectively.
The following result shows that while ${X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$ is not necessarily the best linear predictor of $X_{i,t+h}$ based on the present and past of $\{(\chi_{jt}^{\scriptscriptstyle{\text{\rm stat}}},\xi_{it}^{\scriptscriptstyle{\text{\rm stat}}})\}$, ${X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}$ is the best linear predictor of $X_{i,t+h}$ based on the present and past of $\{(\chi_{jt}^{\scriptscriptstyle{\text{\rm dyn}}},\xi_{it}^{\scriptscriptstyle{\text{\rm dyn}}})\}$ and a.s. outperforms ${X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$. Best, here and in the sequel, is to be understood in the classical sense of minimal expected squared prediction error.
Denote by ${\mathcal{H}}^{\bf X}_{t{\rceil}}$, ${\mathcal{H}}^{{\bm{\chi}}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}$, ${\mathcal{H}}^{{\bm{\chi}}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}$, ${\mathcal{H}}^{{\xi}_{i}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}$, ${\mathcal{H}}^{{\xi}_{i}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}$, ${\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}^{{\bm{\chi}}\,{\xi_{i}}}$, and ${\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}\,{\xi_{i}}}$ the Hilbert spaces spannedby $\{X_{is}|\,i\in{\mathbb{N}},s\leq t\}$, $\{\chi_{js}^{\scriptscriptstyle{\text{\rm stat}}}|\,j\in{\mathbb{N}},s\leq t\}$, $\{\chi_{js}^{\scriptscriptstyle\text{\rm dyn}}|\,j\in{\mathbb{N}},s\leq t\}$, $\{\xi_{is}^{\scriptscriptstyle\text{\rm stat}}|\,s\leq t\}$, $\{\xi_{is}^{\scriptscriptstyle\text{\rm dyn}}|\,s\leq t\}$, $\{(\chi_{js}^{\scriptscriptstyle{\text{\rm stat}}},\xi_{is}^{\scriptscriptstyle{\text{\rm stat}}})|\,j\in{\mathbb{N}},s\leq t\}$, and $\{(\chi_{js}^{\scriptscriptstyle{\text{\rm dyn}}},\xi_{is}^{\scriptscriptstyle{\text{\rm dyn}}})|\,j\in{\mathbb{N}},s\leq t\}$, respectively. If componentwise forecasts are adopted for idiosyncratic components, the best linear $h$-step ahead predictors of $X_{t+h|t}$ based on the static approximate factor model and the GDFM decompositions, are the projections
${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}*}\coloneqq\text{\rm Proj}\left({X}_{i,t+h}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}^{{\bm{\chi}}\,{\xi_{i}}}\right)\quad\text{and}\quad{X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}*}\coloneqq\text{\rm Proj}\left({X}_{i,t+h}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}\,{\xi_{i}}}\right),$
respectively. For the common and idiosyncratic components, these best linear predictors are
${\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}\coloneqq\text{\rm Proj}\left({\chi}_{i,t+h}^{\scriptscriptstyle{\text{\rm stat}}}\ |\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}^{{\bm{\chi}}}\right),\qquad\ \ \ \ {\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}\coloneqq\text{\rm Proj}\left({\xi}_{i,t+h}^{\scriptscriptstyle{\text{\rm stat}}}\ |\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}^{{\xi}_{i}}\right),$
${\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}\coloneqq\text{\rm Proj}\left({\chi}_{i,t+h}^{\scriptscriptstyle{\text{\rm dyn}}}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}}\right),\quad\text{and}\quad{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}\coloneqq\text{\rm Proj}\left({\xi}_{i,t+h}^{\scriptscriptstyle{\text{\rm dyn}}}\ |\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\xi}_{i}}\right).$
Now, since ${\chi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$ is orthogonal to all $\xi_{is}^{\scriptscriptstyle\text{\rm dyn}}$’s, hence to ${\mathcal{H}}^{{\xi}_{i}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}$, and since $\xi_{it}^{\scriptscriptstyle\text{\rm dyn}}$ is orthogonal to all ${\chi}_{js}^{\scriptscriptstyle\text{\rm dyn}}$’s, hence to ${\mathcal{H}}^{{\bm{\chi}}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}$,
${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}*}$ $=$ $\text{\rm Proj}\left({\chi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}\,{\xi_{i}}}\right)+\text{\rm Proj}\left({\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}\,{\xi_{i}}}\right)$
$=$ $\text{\rm Proj}\left({\chi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}}\right)+\text{\rm Proj}\left({\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}|\,{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\xi}_{i}}\right)={\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}+{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}={X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}}.$
In the GDFM approach, adding (as in (9.2)) the optimal predictor ${\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}$ of ${\chi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$ and the optimal marginal predictor ${\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}$ of ${\xi}_{i,t+h}^{\scriptscriptstyle\text{\rm dyn}}$, thus, yields the best linear predictor ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}*}$: this additive combination of common and idiosyncratic forecasts, therefore, is fully justified. This is not the case in a static approximate approach: ${\chi}_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$, indeed, is orthogonal to $\xi_{i,t+h}^{\scriptscriptstyle\text{\rm stat}}$ only, hence fails, in general, to be orthogonal to ${\mathcal{H}}^{{\xi}_{i}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}$. Similarly, $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ is orthogonal to ${\chi}_{jt}^{\scriptscriptstyle\text{\rm stat}}$ for all $j\in{\mathbb{N}}$ but not to ${\mathcal{H}}^{{\bm{\chi}}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}$. As a consequence, adding ${\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$ and ${\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$, as a rule, does not yield the best linear predictor ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}*}$.
Moreover, ${X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}={X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}*}$ outperforms the (hardly implementable) ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}*}$ which, in turn, outperforms the suboptimal ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}}={\chi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}+{\xi}_{i,t+h|t}^{\scriptscriptstyle\text{\rm stat}}$. Indeed, ${\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}\,{\xi_{i}}}$, which contains ${\chi}_{js}^{\scriptscriptstyle\text{\rm weak}}$ for all $j\in{\mathbb{N}}$ and $s\leq t$, includes ${\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}^{{\bm{\chi}}\,{\xi_{i}}}$, which only contains ${\chi}_{is}^{\scriptscriptstyle\text{\rm weak}}$, $s\leq t$. Hence, ${\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm dyn}}}^{{\bm{\chi}}\,{\xi_{i}}}\supseteq{\mathcal{H}}_{t{\rceil}\scriptscriptstyle{\text{\rm stat}}}^{{\bm{\chi}}\,{\xi_{i}}}$ and the squared prediction error $\left[X_{i,t+h}-{X}_{i,t+h|t}^{\scriptscriptstyle\text{\rm dyn}}\right]^{2}$ is a.s.  less than or equal to the (generally unimplementable) $\left[X_{i,t+h}-\widehat{X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}*}\right]^{2}\!$, which in turn is a.s.  less than or equal to $\left[X_{i,t+h}-{X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}}\right]^{2}$.
We have thus proved the following result.
Proposition 2.
Let $\bf X$ be second-order stationary and satisfy both (3.4) and (5.3). Then,
  1. (i)
    ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}}$ defined in (9.1) is not, in general, the best $h$-step ahead linear predictor of $X_{i,t+h}$ based on the present and past of $\{(\chi_{jt}^{\scriptscriptstyle{\text{\rm stat}}},\xi_{it}^{\scriptscriptstyle{\text{\rm stat}}})|\,j\in{\mathbb{N}}\}$;
  2. (ii)
    ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}}$ defined in (9.2) is the best $h$-step ahead linear predictor of $X_{i,t+h}$ based on the present and past of $\{(\chi_{jt}^{\scriptscriptstyle{\text{\rm dyn}}},\xi_{it}^{\scriptscriptstyle{\text{\rm dyn}}})|\,j\in{\mathbb{N}}\}$;
  3. (iii)
    ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}}$ outperforms the best $h$-step ahead linear predictor of $X_{i,t+h}$ based on the present and past of $\{(\chi_{jt}^{\scriptscriptstyle{\text{\rm stat}}},\xi_{it}^{\scriptscriptstyle{\text{\rm stat}}})|\,j\in{\mathbb{N}}\}$, hence also outperforms ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}}$.
Similar statements could be made for predictors ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}}$ and ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}}$ based on alternative idiosyncratic prediction strategies (such as sparse VAR prediction): irrespective of that strategy, the key reason explaining the superiority of ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm dyn}}}$ over ${X}_{i,t+h|t}^{{\scriptscriptstyle\text{\rm stat}}}$ is the mutual orthogonality at all leads and lags of the dynamic common and idiosyncratic components. Yet another reason for adopting the GDFM rather than the Chamberlain and Rothschild static approximate factor approach!
@@ 10 Conclusions
    Since (as shown by Gersing (2023) and Gersing et al. (2023)) it is nesting the classical and widespread static model of Chamberlain (1983) and Chamberlain and Rothschild (1983), the General Dynamic Factor Model of Forni et al. (2000) is uniformly preferable; the benefit is the incorporation into the common space of the GRD-weak factors which consist of statically non-pervasive leads and lags of statically strong factors and, in the static approach, are treated as idiosyncratic;
@@
    Gersing-Rust-Deistler-weak factors “are everywhere,” but in the GDFM approach are incorporated into the space of (dynamically) common components, where they can be efficiently exploited—which explains the empirical finding that the dynamic approach outperforms the static one even when the assumptions of the static factor model are satisfied;
@@
    in forecasting problems, idiosyncratic components (which are not error terms) should not besystematically discarded, as they may be large and have high predictive power; unless orthogonality at all leads and lags is imposed between the statically common and idiosyncratic components (which is a very strong assumption), efficiently combining their predictors into predictors for $\bf X$, however, is not obvious while, in sharp contrast, summing dynamically common and dynamically idiosyncratic predictors (which are mutually orthogonal at all leads and lags) yields optimal linear predictors for $\bf X$.
These conclusions about the advantages of the GDFM over the approximate static model of Chamberlain and Rothschild are likely to extend to spatio-temporal data (for which a GDFM representation result exists, see Barigozzi et al. (2024)), to locally stationary and piecewise stationary high-dimensional time series (for which a GDFM approach is available, see Barigozzi et al. (2021) and, for the piecewise stationary case, Cho et al. (2024), without representation result, though), to functional ones (see Hallin et al. (2023), where a static representation result is established), and to high-dimensional tensor-valued time series (for which a dynamic approach similar to the GDFM still needs to be developed).
Appendix: Proof of Proposition 1
The classical Weyl inequalities straightforwardly entail that (B) implies (A). The proof of Proposition 1, thus, essentially consists of establishing that (A)—that is, the assumption (5.3) on the eigenvalues $\lambda^{(n)}_{X;j}$ of the covariance matrices ${\bm{\Gamma}}^{(n)}_{X}\coloneqq{\rm E}[{\bf X}^{(n)}_{t}{\bf X}^{(n)\prime}_{t}]$ of ${\bf X}^{(n)}_{t}\coloneqq(X_{1t},\ldots,X_{nt})^{\prime}$ for $n\in{\mathbb{N}}$—implies that the conditions (a)–(f) on the static approximate factor model decomposition (2.1) of Chamberlain and Rothschild (along with (d2)) are satisfied by the canonical decomposition (5.2). Essentially, this consists in showing that, under (A), the covariance matrices ${\bm{\Gamma}}^{(n)}_{\chi}\coloneqq{\rm E}[{\bm{\chi}}^{(n)}_{t}{\bm{\chi}}^{(n)\prime}_{t}]$ and ${\bm{\Gamma}}^{(n)}_{\xi}\coloneqq{\rm E}[{\bm{\xi}}^{(n)}_{t}{\bm{\xi}}^{(n)\prime}_{t}]$ of the statically common and idiosyncratic components ${\bm{\chi}}^{(n)}_{t}\coloneqq(\chi_{1t}^{\scriptscriptstyle\text{\rm stat}},\ldots,\chi_{nt}^{\scriptscriptstyle\text{\rm stat}})^{\prime}$ and ${\bm{\xi}}^{(n)}_{t}\coloneqq(\xi_{1t}^{\scriptscriptstyle\text{\rm stat}},\ldots,\xi_{nt}^{\scriptscriptstyle\text{\rm stat}})^{\prime}$ defined in (5.2) and their eigenvalues are such that
@@
    the $r$ non-zero eigenvalues $\lambda^{(n)}_{\chi;1},\ldots,\lambda^{(n)}_{\chi;r}$ of ${\bm{\Gamma}}^{(n)}_{\chi}$ tend to infinity as $n\to\infty$ (it readily follows from (i) that $\lambda^{(n)}_{\chi;j}=0$ for all $r+1\leq j\leq n$);
@@
This takes care, for (5.2), of conditions (b)–(e), while conditions (a) and (f) are automatically satisfied in view of the definition of $\chi_{it}^{\scriptscriptstyle\text{\rm stat}}$ and $\xi_{it}^{\scriptscriptstyle\text{\rm stat}}$ in (5.2). The uniqueness of the decomposition satisfying (a)–(f) then entails that (5.2) and (2.1) coincide.
@@
${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}=\left\{z=\sigma\lim_{n\to\infty}w^{(n)}_{*}{\big{|}}\,(w^{(n)}_{*})_{n\in{\mathbb{N}}}\in{\mathcal{W}}_{r\rceil*},\,\sigma\in{\mathbb{R}}_{+}\right\}.\vspace{-1.5mm}$